OpenStudy (anonymous):

Write an equation in slope-intercept form for a line with a y-intercept of 8 and a slope of -1/4. Answer y = 8x - 1/4 y - 8 = -1/4 x y = -1/4 x + 8 -1/4 x + y = 8

OpenStudy (anonymous):

Use The Law : y=mx+c where m=slope and c=intercepting y at .

OpenStudy (anonymous):

y=-1/4+8?

OpenStudy (anonymous):

Pretty sure you got it. Y=-1/4+8

OpenStudy (anonymous):

U missed the x after the slobe :D y=m(X)+c

OpenStudy (anonymous):

yes i know! haha

OpenStudy (anonymous):

Slope* lol..^_^

OpenStudy (anonymous):

okay i have one more question!

OpenStudy (anonymous):

^ Nice eye.